Fermi National Laboratory
 

Accelerator Update for December 1, 2006 - December 4, 2006

 

Friday December 1

The day shift began with Operations monitoring store 5099 (42.21E30) and a Recycler stash (217.6E10), with the Antiproton Source (Pbar) stacking (54.0E10), with Linac, Booster, the Recycler, and the Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

The Neutron Therapy Facility reported at 8:19 AM that they would treat three patients this morning.

At 9:02 AM, two power supplies (IP7041 & 2) tripped off during the stash. Operations switched to the NuMI only mode and then closed vacuum valves and reset vacuum pumps that had tripped off in the P150 area. Beam resumed to Pbar at 9:13 PM.

At 11:21 AM, the MiniBooNE horn tripped off due to stripline air indication. It reset okay.

The Recycler completed a stash of antiprotons at 12:31 PM.

At 1:53 PM, the NuMI beam permit tripped off with an indication of beam in the kicker gap. The cause was a beam position problem in the AP1 line.

A MI expert began a study at 2:30 PM.

At 2:59 PM, the 750keV line buncher tripped off. Operators and a Linac expert investigated.

At 3:33 PM, the Linac expert reported a short in the buncher anode power supply. He contacted another expert to help.

NuMI experts reported at 4:22 PM that they would be pulsing their spare horn this weekend.

A Linac expert reported at 6:53 PM, that the buncher power supply was operational.

Operations terminated store 5099 around 6:56 PM.

At 6:59 PM, Operations reported that they had established beam to NuMI, MiniBooNE and Pbar, but that the Pbar beam wasn’t making it to the Debuncher.

Operations reported at 7:19 PM that Pbar resumed stacking before the cause of the problem was found.

Operations put the TeV into shot setup at 7:22 PM.

Operations established store 5100 at 9:15 PM with an initial luminosity of 177.8E30.

The Recycler completed a stash at 11:14 PM.

Saturday December 2

The midnight shift began with Operations monitoring store 5100 (101.21E30) and a stash (53.3E10), with Pbar stacking (37.4E10), and with NuMI and MiniBooNE taking beam.

The Recycler completed a stash around 2:07 AM.

At 4:14 AM, store 5100 aborted due to a Tevatron (TeV) power supply trip (QDE2). There was no quench. The supply wouldn’t reset. Operations called in an expert.

Operations turned on a TeV RF station (TRF8) at 5:05 AM after consulting with an expert. It had tripped off on a heater fault.

At 5:36 AM, Operations reported an Accumulator power supply tripped on a ground fault indication (A:IB). Operations called in an expert.

At 5:39 AM, TRF8 tripped off again. Operations contacted and expert and then reset the station.

At 5:45 AM, TRF8 tripped off yet again on a heater fault, only this time it wouldn’t reset. Operations called in an expert.

Operations reported that the TeV power supply QDE2 was repaired.

D0 requested an access to their collision hall. Operators had the area prepared for access by 7:37 AM.

At 7:52 AM, an expert reported that the Accumulator A:IB ground fault problem was in the tunnel. Operators began preparing the tunnel for access.

At 8:53 AM, the Pbar expert reported that the ground fault on A:IB was due to a LCW leak. Operators called in LCW technicians.

TeV experts reported at 9:15 AM that the TRF8 heater was fixed. It was a bad 5-volt power supply.

Operations reported at 11:43 AM that techs had replaced a LCW water hose. The expert had also noticed a problem with a dipole shunt on another device; he replaced the shunt.

Safety system experts reported at 12:05 PM that they had found a bad power supply for the TeV sector E3 electrical safety system that feeds the E2 system that caused the TeV abort.

Pbar resumed stacking at 12:16 PM.

The Recycler completed a stash at 3 PM.

Operations put the TeV into shot setup around 4:32 PM.

Operations established store 5103 at 6:58 PM with an initial luminosity of 105.62E30.

The Recycler completed a stash around 9:17 PM.

The Recycler completed its last stash for Saturday around 11 PM.

Sunday December 3

The midnight shift began with Operations monitoring store 5103 and a stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:07 AM, CDF’s solenoid quenched.

The Recycler completed the second and last stash for the midnight shift around 1:45 AM.

At 2:01 AM, the Booster west anode power supply (WAPS) tripped off for the first of three times in 16 minutes. These trips occur when Booster is in the NuMI/MiniBooNE only mode during cogging. It reset each time.

CDF experts reported at 4:06 AM that they were powering their detector.

At 4:21 AM, MiniBooNE reported that they were having data acquisition problems. Operations turned their beam off. They came back online at 6:45 AM.

At 8:43 AM, the NuMI beam permit tripped due to a BPM problem and wouldn’t reset. Operations turned off their beam and then called an expert while operators continued the investigation.

At 9:11 AM, Operations reported beam position problems in MI. An operator rebooted the low level RF (LLRF) system and that reboot enabled Operations to send beam to Pbar for stacking or to NuMI experimental use, but not beam to both Pbar and NuMI. At 9:38 AM there seemed to be a problem with the Time Line Generator (TLG).

At 10:45 AM, Operations reported a possible corrupted table for the MI LLRF. Operations contacted many experts.

At 11:42 AM, store 5103 disappeared. There was no quench. Operators investigated and then called experts.

Operations reported at 12:20 PM that the MI LLRF problems were resolved. MI experts helped operators rebuild the corrupted table files.

A TeV expert reported at 1:10 PM that he thought the lost store was due to a bad hard wire loop controller at B3. Operations called in power supply experts.

The Recycler completed a stash at 5:21 PM.

Operations put the TeV through dry squeeze at 5:50 PM.

Operations established a 14x0 study-store (14 bunches of protons and zero bunches of antiprotons) at 6:50 PM.

Operations put the TeV into shot setup at 7:16 PM.

Operations established store 5105 at 8:56 PM with an initial luminosity of 187.86E30.

At 8:57 PM, the Booster WAPS tripped off and wouldn’t reset until 9:03 PM.

Monday December 4

The midnight shift began with Operations monitoring store 5105 and stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

The Recycler completed the second and last shot for the midnight shift at 5:43 AM.

The Plans for Monday

The plans for today are to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/06/2006    email Fermilab